OpenStudy (anonymous):

Look at the expression below. 2 + 4xy + 8x + 10 - 3x + 5y a. Rewrite the expression so that like terms are grouped together. b. Simplify the expression I will medal and Fan

OpenStudy (alones.):

2+4xy+8x+10−3x+5y =2+4xy+8x+10+−3x+5y Combine Like Terms: =2+4xy+8x+10+−3x+5y =(4xy)+(8x+−3x)+(5y)+(2+10) =4xy+5x+5y+12

OpenStudy (anonymous):

ohh thx, but which is a and which is b?

OpenStudy (alones.):

It's all together i solved the problem

OpenStudy (anonymous):

ohh lol srry, thx

OpenStudy (alones.):

haha no your fine<3

OpenStudy (anonymous):

k good